Problem 15.125

Water is discharged from a nozzle with a velocity of 8 m/s and strikes the blade mounted on the 18-kg cart. The nozzle has a diameter of 50 mm and the density of water is ρw = 1000 kg/mg3.

A: Determine the tension developed in the cord, needed to hold the cart stationary.

B: Determine the normal reaction of the wheels on the cart.
Answer #1

Answer: Page 1 15-125 ) praw the free body diagram of the control volume. flow rate of the water by calculate the mass Usingpage ② Substitute 1000 kg/m² for Swater, 0.05 m ford, and 8 m/s for v am de = 1000x T X0.05) x8 = 1000 X 0.7853X0.0025 X 8 =page (² -F = 15-706X8( cosuso-1) = 15.706 X 8 (0-7071-1) = 15.706 X8 (-0.2929) Fx = -36.8N Fx = 36.8N force of calculate thepage 4 Draw the free body diagram of the carl- calculate the tension in the cord, and normal reaction of the wheels by usingPages { Fy=0 N-mcg - Fy=0 flere me is the mass of the cart, and g is the acceleration due to gravity. substitute 18kg for me,